OpenStudy (xapproachesinfinity):

I want to prove this directly \[2^{\frac{1}{n+1}}

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

you could try showing that \[\frac{2^{\frac{1}{n+1}}}{n+1}<1\]

OpenStudy (xapproachesinfinity):

i actually tried something similar with logs

OpenStudy (misty1212):

besides the fact that it is completely obvious right?

OpenStudy (misty1212):

the limit is zero for sure

OpenStudy (xapproachesinfinity):

well it is, but that's not enough

OpenStudy (xapproachesinfinity):

analytically yes!

OpenStudy (misty1212):

you want something more complicated? it is true if \(n=1\) now show the function is decreasing

OpenStudy (xapproachesinfinity):

i showed it with calculus i don't want to use calculus somehow haha i showed it is decreasing

OpenStudy (xapproachesinfinity):

direct proof i meant no calculus assumed. just elementary tools :)

OpenStudy (rational):

\[2^{\frac{1}{n+1}}<n+1~~for~~n\ge 1\] is same as showing \[2<2^n~~for~~n\ge 2\]

OpenStudy (thomas5267):

\[ 2^{\frac{1}{n+1}}=2^{\frac{n+2}{(n+1)(n+2)}}>2^\frac{n+1}{(n+1)(n+2)}=2^{\frac{1}{n+2}}\text{ for }n\geq1\\ 2^{\frac{1}{n+1}}>2^{\frac{1}{n+2}}>2^{\frac{1}{n+3}}>2^{\frac{1}{n+4}}>\cdots \\ \text{But:}\\ n+1<(n+1)+1<(n+2)+1<\cdots\\ \text{So:}\\ \cdots<2^\frac{1}{5}<2^\frac{1}{4}<2^\frac{1}{3}<\underbrace{2^\frac{1}{2}<2}_{n=1}<3<4<\cdots \]